# Host Galaxy Demographics Of Individually Detectable Supermassive Black-hole Binaries with Pulsar Timing Arrays

### v1

@@ 5 Conclusions
Individually resolvable single sources producing GW are expected to be detectable by PTAs in the near future. The sky-localization of PTAs is relatively poor and thus, once single sources are detected, it will be unlikely that the true host galaxy will be uniquely pinpointed. Various filtering and selection operations based on consistency between the GW posterior and galaxy catalog values will help reduce the number of plausible hosts [79], but many possibilities may still remain. It is thus critical to analyze the properties of the host galaxies of the most detectable SMBHBs to use as a guide when searching for EM counterparts.
We used the Illustris cosmological simulations coupled with semi-analytic post-processing to identify host galaxies at the time SMBHBs become PTA-detectable. We identified host galaxy properties weighted by the binary detection probability (DP), and compared them to the overall galaxy population. Using the DP, we found that there was an average of 0.14 systems per realization that have a DP greater than $90\%$. Of the various host galaxy properties, the stellar metallicity has the most contrast between hosts of detectable binaries and that of all galaxies. The original unweighted population peaks at $\sim$0.6 solar metallicities while the DP weighted distribution peaks at twice solar metallicity. The color–magnitude diagram gives an indication of the visual appearance of the most detectable galaxies. They are far more red and far brighter, indicating they are most likely large elliptical galaxies with significantly diminished star formation.
Each of our distributions of host galaxy properties gives us a tool to rank prospective gaalxies by their likeliness of harboring a PTA-detected single source; this is very useful information to aid the goal of multi-messenger astronomy of SMBHBs. A natural next step for future work is to investigate the properties of SMBHBs that would produce detectable electromagnetic counterpart signatures through mechanisms like shifted broad-line regions, accretion variability, Dopper boosting, and self-lensing. Side by side with the population that is detectable in GWs by PTAs—as we have investigated in this work—one could find the overlap in the two populations, and thus examine the typical properties of binaries (and their hosts) that are the most likely to be multi-messenger detectable.
@@ Acknowledgments
We thank Maria Charisi for extensive discussions that improved the quality of this work. KC acknowledges support from the Vanderbilt University Data Science Institute Summer Research Program (DSI-SRP). SRT acknowledges support from NSF AST-2007993, AST-2307719, and an NSF CAREER PHY-2146016, and a Vanderbilt University College of Arts & Science Dean’s Faculty Fellowship. SRT and LK are members of the NANOGrav collaboration, which receives support from NSF Physics Frontiers Center award number 1430284 and 2020265. This work was conducted in part using the resources of the Advanced Computing Center for Research and Education (ACCRE) at Vanderbilt University, Nashville, TN.

### v2

@@ 5 Conclusions
Individually resolvable single sources producing GW are expected to be detectable by PTAs in the near future. The sky-localization of PTAs is relatively poor and thus, once single sources are detected, it will be unlikely that the true host galaxy will be uniquely pinpointed. Various filtering and selection operations based on consistency between the GW posterior and galaxy catalog values will help reduce the number of plausible hosts [91], but many possibilities may still remain. It is thus critical to analyze the properties of the host galaxies of the most detectable MBHBs to use as a guide when searching for EM counterparts.
We used the Illustris cosmological simulations coupled with semi-analytic post-processing to identify host galaxies at the time MBHBs become PTA-detectable. We identified host galaxy properties weighted by the binary detection probability (DP), and compared them to the overall galaxy population. Using the DP, we found that there was an average of 0.14 systems per realization that have a DP greater than $90\%$. Of the various host galaxy properties, the stellar metallicity has the most contrast between hosts of detectable binaries and that of all galaxies. The original unweighted population peaks at $\sim$0.6 solar metallicities while the DP weighted distribution peaks at twice solar metallicity. The color–magnitude diagram gives an indication of the visual appearance of the most detectable galaxies. They are far more red and far brighter, indicating they are most likely large elliptical galaxies with significantly diminished star formation.
Each of our distributions of host galaxy properties gives us a tool to rank prospective galaxies by their likeliness of harboring a PTA-detected single source; this is very useful information to aid the goal of multi-messenger astronomy of MBHBs. This demographic profile of binary hosts could be combined with other recent work showing that galaxy morphology [12] and stellar kinematics [11] could signpost the presence of a massive central binary. A natural next step for future work is to investigate the properties of MBHBs that would produce detectable electromagnetic counterpart signatures [e.g., 87, in an SKA context]. For a review of these counterparts, we point the reader to D’Orazio & Charisi [28]. Some of the most promising signatures are shifted broad-line regions [e.g. 60], accretion variability [e.g. 40], Dopper boosting [e.g. 32], and self-lensing [e.g. 30, 68]. However, due to the extreme intrinsic variability of AGN, each of these signatures is subject to substantial contamination from false positives. Side by side with the population that is detectable in GWs by PTAs—as we have investigated in this work—we can thus drastically improve the chances of identifying multi-messenger sources in the next few years.
@@ Acknowledgments
We thank the anonymous referees for their comments. We thank Maria Charisi for extensive discussions that improved the quality of this work. KC acknowledges support from the Vanderbilt University Data Science Institute Summer Research Program (DSI-SRP). SRT acknowledges support from NSF AST-2007993, AST-2307719, and an NSF CAREER PHY-2146016, and a Vanderbilt University College of Arts & Science Dean’s Faculty Fellowship. SRT and LK are members of the NANOGrav collaboration, which receives support from NSF Physics Frontiers Center award number 1430284 and 2020265. This work was conducted in part using the resources of the Advanced Computing Center for Research and Education (ACCRE) at Vanderbilt University, Nashville, TN.